OpenStudy (anonymous):

I'm trying to solve by elimination but im confused!? Please help! 2x + 5y = −3 −15x − 5y = −10

OpenStudy (anonymous):

@experimentX

OpenStudy (experimentx):

add both sides of both

OpenStudy (anonymous):

so add 3 to the top and 10 to the bottom

OpenStudy (experimentx):

(2x + 5y)+(−15x − 5y) = −3+(-10) rest is easy

OpenStudy (anonymous):

1 second let me try

OpenStudy (anonymous):

I got 17x=13

OpenStudy (anonymous):

hello

OpenStudy (anonymous):

i'll help you lol

OpenStudy (experimentx):

no ... do it again, 17x on left side is incorrect

OpenStudy (anonymous):

x = 1

OpenStudy (anonymous):

it's -13x=-13

OpenStudy (anonymous):

the equation was literally set up for elimination

OpenStudy (anonymous):

just add top and bottom

OpenStudy (anonymous):

oh so I did it wrong

OpenStudy (anonymous):

no you just messed up

OpenStudy (anonymous):

-3-10= -13

OpenStudy (anonymous):

2x-15x= -13x

OpenStudy (anonymous):

-13x =-13

OpenStudy (anonymous):

thank bro you're a life saver!